OpenStudy (anonymous):

4/11 divided by 4/9 =?

OpenStudy (anonymous):

9/11

OpenStudy (anonymous):

4/11 divided by 4/9= 9/11 This site might help http://www.mathsisfun.com/fractions_division.html

OpenStudy (anonymous):

4/11 ÷ 4/9 = (4 × 9) / (4 × 11) = 36/44 = 9/11

OpenStudy (anonymous):

anymore

OpenStudy (anonymous):

i said that